### v2

# A Parametric Study on the Pulsation of Burner-Stabilized CH4-O2 Flames Moderated by CO2 Addition – Frequencies, Modes and Regime Diagrams

Abstract
This study investigated the pulsating instability of burner-stabilized premixed CH4-O2 flames at various levels of CO2 dilution. Experiments were conducted using a water-cooled porous-plug burner of 18 mm diameter over a wide range of mixture compositions and flow rates, during which time-resolved measurements of flame chemiluminescence and gas temperature were obtained. The primary oscillation frequencies of the pulsating flames were determined using fast Fourier transform and harmonic power analysis. Phase-locked analysis of the chemiluminescence images revealed an interesting mode-transition phenomenon of the flame oscillations. Under fuel-rich conditions with relatively low heat release rates and low flow rates, the flames exhibited quasi-periodic single-mode oscillations. These oscillations were modulated by low-frequency flame flickering instabilities at elevated flow rates, which created sidebands around the primary oscillation frequency. At higher heat release rates, the flickering instability further triggered mode splitting, eventually leading to multi-mode oscillations. Regime diagrams of the flame oscillation modes, as well as the stability boundaries, were obtained under various fuel flow rates. These findings can be useful for both fundamental research on flame dynamics and practical applications of CO2-moderated oxy-combustion.
keywords:
Pulsating Instability; Oxy-Combustion; Porous-Plug Burner; Chemiluminescence; Laser Absorption Spectroscopy; Phase-Locked Analysis
journal: Fuel
[…]
organization=SKLTCS, CAPT, School of Mechanics and Engineering Science, Peking University, addressline=5 Yiheyuan Road, Haidian District, city=Beijing, postcode=100871, country=China
{highlights}
Novelty and Significance
This work presents, to the authors’ knowledge, the first systematic measurements of pulsating flame instabilities in burner-stabilized premixed methane-oxygen flames across a wide range of experimental conditions (equivalence ratios, dilution factors and flow rates). The combination of a porous-plug burner and oxygen-enriched combustion amplifies weak phenomena that commonly occur under extreme conditions of very high Lewis number, making them observable under moderate conditions that are easily accessible in a laboratory. Additionally, this work reports direct measurements of the spatiotemporal evolution of flame chemiluminescence and gas temperature during pulsation cycles. Based on phase-locked analysis of the data, this study has revealed an interesting mode-transition phenomenon of the flame pulsation. The results of this work should prove useful for both fundamental research on flame dynamics and practical applications of CO2-moderated oxy-combustion.
Author Contributions
Xiangyu Nie: Data curation, Investigation, Writing - original draft. Shuoxun Zhang: Data curation, Investigation, Writing - original draft. Shengkai Wang: Conceptualization, Formal Analysis, Funding acquisition, Methodology, Supervision, Writing — original draft, Writing - review & editing.
1 Introduction
Combustion instabilities, caused either by the interaction of reacting flow with combustor components or by intrinsic processes related to flame dynamics, pose critical challenges to the modern design of power generation and propulsion systems aimed at achieving cleaner and more efficient operations at higher safety levels [1, 2, 3], especially amid the increasing adoption of renewable fuels [4] and carbon-neutral combustion strategies [5]. For example, fuels obtained from renewable sources typically contain a significant portion of lightweight molecules that exhibit kinetic, thermodynamic, and transport properties drastically different from those of conventional large hydrocarbon fossil fuels. These properties can lead to intrinsic flame instabilities within the combustion process itself, even in the absence of surrounding components [6].
An important type of such instability is the thermo–diffusive instability that arises from the imbalance between heat and mass transfer [7, 8, 9]. Depending on the effective Lewis number (Le) of the reactive mixture, the flame can exhibit either spatial cellular structures (at Le $<$ 1) or temporal pulsation of the reaction front (at Le $>$ 1). Although the former has been extensively investigated, especially in lean combustion systems of lightweight fuels such as hydrogen and methane [10, 11, 12, 13, 14, 15, 16, 17, 18], experimental studies on the latter remain relatively scarce.
From a physics perspective, this scarcity is likely due to the criterion for the onset of pulsating instabilities that rendered experimental observation difficult. According to the classic theoretical analysis of Sivashinsky [19], pulsating instability in a freely propagating premixed flame can occur only when $\rm Ze(Le-1)>4(1+\sqrt{3})$, where Ze is the Zeldovich number representing the non-dimensional overall activation energy of the mixture. This requires the Lewis number to be impractically high for most fuel-air mixtures [20], except for very rich hydrogen/air mixtures [21]. Though not being an issue for theoretical and computational studies [22, 23], where arbitrarily large Lewis numbers can be accessed [24, 25, 26], it has posed substantial challenge for general experimental measurements of pulsating instability in freely propagating flames.
In light of this challenge, researchers have resorted to alternative platforms, such as the porous-plug burner, for the experimental investigation of flame pulsation dynamics (for example, see [27, 28]). This type of burner features a flat surface made of porous metal plug, which generates a uniform velocity profile and also acts as a heat sink that helps to anchor the flame. The presence of a porous plug amplifies the flame instability through interactions with the boundary conditions of the burner, thereby making the pulsation phenomena observable over a wider range of conditions [29], especially at much lower Lewis numbers that are more relevant to practical engines and combustors. Furthermore, from a pragmatic perspective, this configuration perhaps better represents realistic situations, as all practical combustors have boundaries.
Some pioneering investigations on the thermo-diffusive pulsation of burner-stabilized flames date back to the 1980s and 1990s. For example, Blackshear et al. [30] observed flame front oscillations in experiments with lean propane and rich methane–air mixtures; Pearlman and Ronney [31] studied a variety of pulsation modes (including concentric rings and rotating spiral waves) in methane, propane, and butane flames in air. Recently, there has been a proliferation of new experimental and computational studies in this area (for example, see [21, 32, 20, 33, 34, 35, 36]), focusing primarily on methane-air and hydrogen-air flames. However, for flames with enriched or pure oxygen as the oxidizer, a comprehensive characterization of the pulsating instability remains lacking. These flames are expected to exhibit substantially different instability boundaries and oscillation behaviors due to their higher volumetric heat release rates.
Extended study on this topic is important not only for advancing fundamental flame dynamics research but also for practical applications of oxy-fuel combustion. One prominent application is in rocket engines powered by liquid oxygen and hydrocarbon fuels [37]. The successful operation of advanced rocket thrust chambers demands robust mitigation of combustion instabilities, particularly inside the gas generator, where combustion of the methane-oxygen mixture occurs at fuel-rich conditions and can be subject to thermo-diffusive pulsation [38]. Aside from propulsion applications, oxy-combustion is also receiving increased attention in power generation as a promising decarbonization strategy that simplifies CO2 capture by eliminating nitrogen from the exhaust stream [39, 40]. In such systems, flue gas recirculation is often required to moderate flame temperature [41, 42], making it essential to account for the effects of CO2 dilution as well. However, as previous studies have suggested [5, 43], the flame instability characteristics of O2/CO2 combustion systems are quite different from combustion in air, because (a) the oxygen enrichment modifies the Zeldovich number by changing the effective active energy and raising the adiabatic flame temperature, and (b) the introduction of CO2 molecules in the unburnt gas mixture modifies its effective Lewis number. These instability characteristics merit further investigation.
The current study aims to extend previous knowledge on pulsating flame instability by systematically exploring the stability boundary, dominant frequency, and modal characteristics of this instability under conditions of oxy-combustion and more realistic thermal boundaries with finite heat loss. The rest of this paper is organized as follows: Section 2 elaborates on the methods of experimental measurement and data analysis; Section 3 presents the measurement results and discusses the frequencies, modes, and regime diagrams of the observed flame oscillation phenomena; and Section 4 provides the concluding remarks.
2 Methods
2.1 Experimental method
The current flame experiments were conducted on a custom-built McKenna-type burner, as shown in Fig. 1. The burner had a circular sintered bronze plug of 18 mm diameter that was water-cooled to maintain a stable temperature. The temperature of the porous plug was continuously monitored using an embedded K-type thermocouple (Omega TJ36-CAXL-020-12). High-purity CH4 (99.99%-grade), O2 (99.999%-grade), and CO2 (99.999%-grade) were supplied to the bottom of the burner, where they were thoroughly mixed by an in-line static mixer before passing through the porous plug. The flow rates of fuel, $\rm O_{2}$, and $\rm CO_{2}$ were precisely controlled by three Alicat MC series mass flow controllers, with typical uncertainties of 0.1%, 0.2%, and 0.5%, respectively.
Multiple diagnostic methods were employed to investigate the flame oscillation dynamics, including: (1) a spatially and spectrally modulated multi-color laser absorption system for spatiotemporally resolved measurements of gas temperature and CO2 mole fraction; (2) an image-intensified high-speed camera for spatiotemporally resolved imaging of the line-of-sight integrated OH* chemiluminescence (as a proxy for the local heat release rate); and (3) a photomultiplier tube (PMT) for microsecond-resolved measurements of global chemiluminescence variations and transient phenomena.
Figure 1: The current experimental setup: (a) schematic of the overall diagnostic system and burner; (b) detailed configuration of the porous-plug burner; (c) example of a burner-stabilized flame; (d) detailed view of the water cooling channel; (e) photograph of the static mixer.
The current laser absorption diagnostic used two distributed-feedback interband-cascade lasers (Nanoplus) to probe 12 absorption features of hot CO2 molecules near the $\nu_{3}$ bandhead at a scan rate of 10 kHz. These transitions have drastically different temperature sensitivities. From the relative intensities of these transitions, the gas temperatures were determined with high sensitivity and robustness. Spatially resolved measurements were enabled by a 2D high-speed parallel beam scanning system consisting of a 2D galvo scanner (Thorlabs, GVS102 2D) and a pair of off-axis parabolic mirrors placed upstream and downstream of the measurement region. The galvo scanner was located at the focus of the upstream off-axis parabolic mirror, ensuring that the reflected beams remained parallel to its axis. The downstream mirror focused the parallel beams on a photodetector that was bandpass-filtered to block the background infrared emission of the flames. A planar measurement speed of 200 Hz was routinely achieved. Further details of this diagnostic method have been documented in a separate work of the authors [44] and are omitted here for brevity.
Time-synchronized measurements of flame chemiluminescence were conducted using an image-intensified high-speed monochromatic camera (Phantom v611) at a frame rate of 10 kHz and a bit depth of 12 bits. To avoid saturation effects, the gain of the image intensifier (Intelligent Scientific Systems, Model EyeiTS) was adjusted to yield a maximum pixel value of approximately 75% of the camera’s full range. For top-view measurements of the flames, a flat mirror was placed 20 cm above the burner. This imaging setup provided a depth of field of approximately 6 mm, effectively restricting the recorded emission to a narrow region surrounding the flame front while blurring the out-of-focus signal from further downstream. This allowed for selective visualization of the reaction zone and minimized interference from background emission. Further details of the imaging setup were documented in a previous study by the authors [45]. The flames were recorded at a pixel resolution of 256 x 256 and a physical resolution of 0.11 mm x 0.11 mm per pixel. During each measurement, a sequence of 10,000 consecutive images was recorded.
[…]
2.2 Frequency analysis and phase-locked cycle averaging of chemiluminescence signal
For a pulsating flame, the frequency content of the total chemiluminescence signal was first analyzed using fast Fourier transform (FFT). The results were expressed in terms of normalized power spectral density (NPSD), as shown in Eqn. (1).
$S(f_{k})=\frac{1}{N\Delta f}\cdot\frac{|\sum_{i=1}^{N}I(t_{i})e^{-2\pi jf_{k}t_{i}}|^{2}}{\sum_{i=1}^{N}I(t_{i})^{2}}$ (1)
In Eqn. (1), $f_{k}=k\Delta f$ denotes the k-th discrete Fourier frequency, with $\Delta f$ being the reciprocal of the total measurement time $T=N\Delta t$ (1 s in the current study). $S(f_{k})$ denotes the NPSD of the chemiluminescence signal evaluated at $f_{k}$. $I(t_{i})$ represents the discrete-time chemiluminescence signal sampled at $t_{i}=i\Delta t$, and $j$ denotes the imaginary unit. $S(f_{k})$ is also expressed in logarithmic units (dB/Hz) as follows.
$\hat{S}(f_{k})=10\log_{10}S(f_{k})$ (2)
A harmonic power analysis method was employed to extract the fundamental frequency of oscillation, $f_{osc}$, by exploiting the fact that the frequency content of a periodic physical event consists of harmonics that are phase-locked, whereas the frequency content of noise or random fluctuations is uncorrelated. Specifically, $f_{osc}$ was determined by solving a constrained optimization problem, namely maximizing the total power of harmonics within a search window $[f_{min},f_{max}]$ estimated from the NPSD result, as shown below.
$f_{osc}=\underset{[f_{min},f_{max}]}{\operatorname{argmax}}\sum_{k=1}^{M}\left|\sum_{i=1}^{N}I(t_{i})e^{-2\pi jkft_{i}}\right|^{2}$ (3)
This optimization problem was solved numerically using a brute-force search algorithm, and to improve computational efficiency and robustness, the number of harmonics was truncated to M = 5. In the current study, this method was able to determine the primary oscillation frequency with a typical uncertainty of less than 0.1 Hz (defined by the frequency offset at 95% maximum total power), which was an order of magnitude better than the conventional FFT.
Once the primary oscillation frequency was determined, the chemiluminescence images of the pulsating flames were further analyzed using phase-locked cycle averaging. The recorded image sequence was divided into sub-sequences corresponding to individual cycles of oscillation, and images captured at the same relative time within each cycle were averaged across different sub-sequences. The starting frame of each sub-sequence was fine-tuned through cross-correlation analysis, using the spatially integrated chemiluminescence signal as a reference, which effectively removed random jitters in the oscillation cycles. This procedure was executed iteratively until convergence was achieved, usually within 2-3 iterations. This phase-locked cycle averaging successfully removed random fluctuations and revealed the underlying physical trends consistent across cycles, without degrading the effective temporal resolution of the current measurements (100 $\mu s$).
2.3 Axisymmetric reconstruction of temperature fields from laser absorption measurement
Spatially resolved measurements of local gas temperature were achieved by scanning the laser beams across the flame. An example measurement of a pulsating CH4-O2 flame at lower flow rates ($\dot{m}_{\rm CH_{4}}$ = 0.20 SLPM and $\dot{m}_{\rm O_{2}}$ = 0.30 SLPM, respectively) is illustrated in Fig. 2. Under the specific conditions of this measurement, the flame exhibited a regular mode of oscillation that was periodic in time and nearly axisymmetric in space, as confirmed by time-synchronized chemiluminescence imaging. For oscillating flames exhibiting axial symmetry, a reconstruction scheme using constrained optimization was developed to infer the spatial distribution of temperature from the absorbance data at different beam locations.
Figure 2: Example measurement of a premixed CH4-O2 flame pulsating in a single axisymmetric mode under the conditions of $\dot{m}_{\rm CH_{4}}$ = 0.20 SLPM and $\dot{m}_{\rm O_{2}}$ = 0.30 SLPM. (a) A spectrogram of the absorbance signal at different beam locations, with data obtained from various oscillation cycles at a fixed relative time of 10 ms. (b) The integrated absorbance of an isolated feature at different beam locations. (c) The Spatial distribution of gas temperature reconstructed from integrated absorbances under the assumption of axial symmetry.
The line-of-sight absorbance $\alpha(t;y_{i})$ was determined at each beam location $y_{i}$ from the transmitted laser intensity $I(t;y_{i})$ and a reference intensity profile $I_{0}(t)$ measured in the absence of absorption, based on the Beer-Lambert relation:
$\alpha(t;y_{i})$ $=-\ln\big{[}I(t;y_{i})/I_{0}(t)\big{]}$ (4)
From the time sequence of $\alpha(t;y_{i})$, an absorbance spectrum $\alpha(\nu;y_{i})$ was obtained based on the laser wavenumber modulation profile $\nu(t)$ characterized prior to the flame experiment. A total of 12 absorption features in the spectral region of 2392.5 to 2397 $\rm cm^{-1}$ were analyzed, and the integrated absorbance over each selected transition feature (denoted by $[\nu_{k-},\nu_{k+}]$) was used in the axisymmetric reconstruction, i.e.,
[…]
To determine the spatial distribution of gas temperature, the entire spatial domain of interest was discretized using a number of sample points $r_{s}$ along the radial direction. These sample points were selected based on the effective spatial resolution of the current laser diagnostic (1 mm, as defined by the laser beam diameter). At each sample point, a local gas temperature $T_{s}$ and a CO2 mole fraction $X_{s}$ were assigned. The integrated absorbance of each absorption feature was evaluated first at each sample point $A^{0}_{k}(r_{s};T_{s},X_{s})$ and then summed along each line of sight (see Fig. 2(b)), yielding:
$A_{k}^{0}(r_{s};T_{s},X_{s})$ $=PX_{s}S_{k}^{0}(T_{s})$ (6)
and
[…]
In the above equations, a constant flame pressure of $P$ = 1 atm was assumed, and the linestrength of each transition feature, $S_{k}^{0}$, was calculated as a function of the local gas temperature from the HITEMP spectroscopy model [46]. The line-of-sight integration was numerically evaluated using Gauss–Legendre quadrature, with linear interpolation between sample points. This approach effectively approximated the integral as a weighted sum of the integrated absorbances at selected sample points. The weight factor $w_{is}$ was a geometric constant independent of the laser wavenumber $\nu$ and the local gas properties $T_{s}$ and $X_{s}$. The integrated absorbances were assembled into two column vectors of $\boldsymbol{A^{0}(T,X)}=\{A^{0}_{k}(r_{s};T_{s},X_{s})\}$ and $\boldsymbol{A}=\{A_{k}(y_{i})\}$, with $\boldsymbol{T}$ and $\boldsymbol{X}$ representing $\{T_{i}\}$ and $\{X_{i}\}$, respectively. All weight factors $\{w_{is}\}$ were compiled into a constant matrix $\boldsymbol{W}$. The optimal values of $\boldsymbol{T}$ and $\boldsymbol{X}$ were determined by solving the following nonlinear optimization problem:
$(\boldsymbol{T^{*},X^{*}})={\operatorname{argmin}}\big{\|}\boldsymbol{A-WA^{0}(T,X)}\big{\|}_{2}^{2}+\lambda_{T}^{2}\big{\|}\boldsymbol{LT}\big{\|}_{2}^{2}+\lambda_{X}^{2}\big{\|}\boldsymbol{LX}\big{\|}_{2}^{2}$ (8)
The last two terms in Eqn. 8 represent smoothness regularization, with $\boldsymbol{L}$ being the discrete Laplacian matrix and $\lambda_{T}$ and $\lambda_{X}$ serving as the regularization strength factors. The values of $\lambda_{T}$ and $\lambda_{X}$ were generally kept small to avoid blurring the reconstructed distributions, while being sufficiently large to ensure numerical stability and suppress random fluctuations; in the current study, they were set to $1\times 10^{-5}$ and $1\times 10^{-1}$, respectively. Eqn. 8 was solved iteratively using the sequential quadratic programming method. The resulting spatial distributions of temperature and CO2 mole fraction are shown in Fig. 4(c), with a typical measurement precision of 1%. Further discussion on the uncertainty analysis, as well as the choice of $\lambda_{T}$ and $\lambda_{X}$, can be found in a recent work by the authors [44].
3 Results and Discussion
A total of 838 pulsating flame experiments were conducted at equivalence ratios ($\phi$) ranging from 1.18 to 2.35 and dilution factors ($\eta$, defined as $\dot{m}_{CH_{4}}/(\dot{m}_{CH_{4}}+\dot{m}_{CO_{2}})$) ranging from 0.26 to 1.00. The primary frequencies, oscillation modes, regime diagrams, and fuel dependence of these pulsating flames are discussed as follows. A summary of the frequencies and modes of these oscillating flames can be found in Supplementary Material I.
3.1 Pulsation frequencies
The primary oscillation frequency ($f_{osc}$) of a pulsating flame was determined using the method described in Section 2.2. Across the range of the current experimental conditions, this frequency varied from 26.4 to 1866.0 Hz, exhibiting a negative dependence on the equivalence ratio and a positive dependence on the dilution factor.
Figure 3: (a) Representative OH* NPSD spectra of pulsating flames at three different conditions. Top (blue): at $\phi$ = 1.72, $\eta$ = 1.00, the NPSD spectrum consists of equally spaced harmonics of a single fundamental frequency, indicating the presence of single-mode oscillation. Middle (green): at $\phi$ = 1.72, $\eta$ = 0.71, low-frequency flickering instability appears, which couples with the single-mode oscillation and creates sidebands around the harmonics. Bottom (red): at $\phi$ = 1.47, $\eta$ = 0.71, multiple fundamental oscillation frequencies (and their harmonics) are observed. $f_{\text{OSC}}$: the primary oscillation frequency; $f_{\text{FI}}$: the flickering instability frequency. (b) Total harmonic power analysis of the three pulsating flames.
Some representative OH* NPSD spectra of pulsating flames are illustrated in Fig. 3. Depending on the specific flow rates of fuel, O2, and CO2, the flame oscillations can exhibit distinctly different characteristics. The top panel of the figure presents a typical example of fuel-rich, undiluted (or weakly diluted), low-speed flames, where both the heat release rate and the mean flow velocity are relatively low. The NPSD spectrum contains equally spaced harmonics of a single fundamental frequency around 84 Hz, with no other frequency content observed. These characteristics indicate that the flame oscillates in a single periodic mode, as corroborated by the single peak in the total harmonic power spectrum displayed in the right panel.
At higher flow velocities, a second, low-frequency oscillation appeared around 10 Hz and created sidebands around the primary oscillation frequency, as shown in the middle panel of the figure. Across the experimental conditions explored in the current study, the frequency of this second oscillation was found to be relatively stable (9 - 12 Hz) and largely independent of the equivalence ratio and dilution level, suggesting that it must have originated from a mechanism different from the high-frequency thermo-diffusive pulsation.
To further investigate the nature of this low-frequency oscillation, additional experiments were conducted using a CMOS camera to monitor the chemiluminescence signal of an undiluted, moderately fuel-rich flame from the side, as illustrated in Fig. 4. Under these conditions, the flame did not exhibit any high-frequency pulsation; only a low-frequency oscillation at a fundamental frequency of 11 Hz was observed. The results indicate that the low-frequency component corresponds to flickering-type flame dynamics that is distinctively different from flame front pulsation. For example, throughout a flame flickering cycle, the base of the flame remained largely stationary, whereas the tail of hot burnt gas downstream of the flame exhibited an oscillating motion of elongation and contraction along the axial direction.
It turns out that the current observation of this rich premixed flame agrees well with classic studies on diffusion flames, in terms of both the overall phenomenon and the range of oscillation frequencies (e.g., [47, 48]). According to these studies, such low-frequency oscillations of the flame result from buoyancy-induced flow surrounding the flame. The velocity difference between the hot burnt gases and the cold ambient air creates an unstable shear layer, which rolls up into toroidal vortices and sheds periodically. The current study thus refers to this low-frequency oscillation as the flame flickering instability, with its frequency denoted as $f_{\rm FI}$.
Figure 4: representative example on the low-frequency oscillation of a methane-oxygen flame at $\phi$ = 1.20, $\eta$ = 1.00. (a) the FFT spectrum of CH* normalized power spectral density (NPSD), (b) sequential side-view images of the flame.
At higher heat release rates (i.e., higher fuel flow rates and/or lower equivalence ratios), the sidebands around the primary oscillation frequency of the flame further split into two or more fundamental oscillation frequencies, each accompanied by its own harmonics, as shown in the bottom panel of Fig. 3. In the current study, this phenomenon is termed multi-mode oscillation. In this context, the corresponding primary oscillation frequency is defined as the fundamental frequency of the stronger mode. Three representative cases of different flame oscillation modes have been selected for further analysis, as discussed in the next section.
3.2 Single-mode oscillation
Fig. 5 presents the dynamic evolution of multiple scalar fields in a single-mode pulsating flame under the conditions of $\dot{m}_{\rm CH_{4}}$ = 0.20 SLPM and $\dot{m}_{\rm O_{2}}$ = 0.30 SLPM. The top panel of the figure illustrates the temporal evolution of chemiluminescence signal in a cycle, where phase-locked averaging was applied to remove random fluctuations and increase the signal-to-noise ratio. The resulting data has been normalized to a relative intensity scale of -1 to 1. Excellent agreement is seen between the pixel-integrated signal of chemiluminescence image and the PMT signal. Within each oscillation cycle of approximately 65 ms, the total chemiluminescence is seen to increase during the first 45 ms and decrease during the last 20 ms.
Eight representative frames of the cycle-averaged chemiluminescence images are shown in Fig. 5(b) in pseudocolors, with red and blue representing high and low chemiluminescence signals, respectively. The mean intensity has been subtracted from each pixel to enhance contrast and highlight dynamic changes. From these images, the pulsating flame is observed to oscillate in a propagating ring mode, with concentric waves of local heat release rate perturbation originating from the edge and traveling toward the flame center, where they merge and terminate before the next oscillation cycle.
It is worth noting that, although the measured images were line-of-sight integrated, no significant cancellation effect was expected in the top-view images because the flame front remained close to a geometrically thin sheet throughout the oscillation cycle. Besides, the flame motion occurred primarily in the vertical direction, as revealed by additional side-view measurement of a single-mode pulsating flame shown in Fig. 6.
Fig. 5(c) illustrates the dynamic evolution of temperature distribution at a height above the burner (HAB) of 2.5 mm. A similar ring-type oscillation can be observed in the temperature distribution, but with a phase lag compared to the chemiluminescence signal. At the specific conditions of this experiment, the maximum local temperature is found to be approximately 2000 K, significantly lower than the adiabatic flame temperature ($\approx$ 3000 K) due to heat loss at the water-cooled burner surface.
Fig. 5(d) shows the evolution of line-of-sight-averaged temperature within one oscillation cycle. In the current study, a high level of measurement precision (on the order of $\pm$ 5 K) has been achieved through cycle averaging, thereby allowing small variations to be discerned. The average temperature is seen to oscillate between 1760 and 1790 K, with a phase lag of 100 - 150 degrees relative to the chemiluminescence signal. The overall 1-$\sigma$ uncertainties in these measurements are estimated to be 2 %, based on the uncertainty contributions from the spectral fitting (less than 0.1%), spectroscopy model (less than 1%), and small fluctuations in the laser intensity baseline (less than 1%).
Fig. 5(e) presents the vertical distribution of average temperature along the burner diameter, measured at different heights above the burner (HAB) and at various times during a single oscillation cycle. Its dynamic evolution reveals an interesting phenomenon that cannot be captured by the chemiluminescence imaging alone, namely, the existence of higher and lower average temperature regions and their vertical motion with time, which implies alternating puffs of hot and cool gas traveling in the vertical direction. Further investigation into the 2D/3D structure of the traveling flame fronts, especially in conjunction with velocity field measurement, may warrant future studies.
Figure 5: Spatiotemporally resolved measurements of a single-mode pulsating flame ($\phi$ = 1.33, $\eta$ = 1.00) at different times in the oscillation cycle. (a) Spatially integrated chemiluminescence signals from the high-speed camera (black) and the PMT (blue), normalized and phase-locked averaged over 90 cycles. (b) Phase-locked averaged chemiluminescence images. (c) Dynamic evolution of the gas temperature distribution at HAB = 2.5 mm. (d) Evolution of the average temperature along the burner diameter. Error bars denote the standard deviation of the measurements. (e) Evolution of the vertical distribution of average gas temperature along the burner diameter within a single oscillation cycle.
Figure 6: Side-view chemiluminescence images of a single-mode pulsating flame at $\phi$ = 2.00, $\eta$ = 1.00.
Figure 7: Chemiluminescence signal illustrating the spatiotemporal modes of flame oscillation. Selected frames of the cycle-averaged chemiluminescence images, after the subtraction of mean intensity, are presented in pseudocolors, with red and blue representing high and low signals, respectively. (a) Single-mode oscillation influenced by flickering instabilities at $\phi$ = 1.72, $\eta$ = 0.71. The dominant flame oscillation mode appears to be a fast centripetal traveling wave that is nearly axisymmetric. The flickering instabilities are manifested as a slow edge-mode perturbation that modulates the relative phase of the fast flame oscillation without significantly affecting its amplitude and spatial distribution. (b) Multi-mode oscillation at $\phi$ = 1.47, $\eta$ = 0.71. The integrated chemiluminescence intensity exhibits a clear beat pattern, indicating the presence of two fundamental oscillation frequencies that interfere with each other. The propagation directions and spatial patterns of flame oscillations during periods of constructive and destructive interference are drastically different.
3.3 Multi-mode oscillation
Two more complex cases of pulsating flames are presented in Fig. 7, including: (a) single-mode oscillations coupled with flickering instabilities, and (b) multi-mode oscillations. In the former case, the dominant flame oscillation mode remains a centripetal traveling wave, but flickering instabilities introduce a slow edge-mode perturbation that modulates its relative phase without significantly affecting its amplitude and spatial distribution. In the latter case, the amplitude of the flame oscillation varies significantly over time.
For example, in a representative measurement shown in Fig. 7(b), the integrated chemiluminescence intensity exhibits a clear beat pattern, indicating the presence of two fundamental oscillation frequencies interfering with each other. During the time interval of constructive interference, the spatial pattern of the flame oscillation remains axisymmetric, while the propagation of the heat release rate perturbation becomes bidirectional – waves originate from a ring zone and travel toward both the center and the edge of the burner. In contrast, during the time interval of destructive interference, the flame oscillation pattern appears drastically different: perturbation waves travel along two separate loops near the top and bottom edges of the burner and meet along the diameter. More complex interactions among three or more modes have also been observed at much higher heat release rates.
Dynamic Mode Decomposition (DMD) analysis was also performed on selected OH* image sequences, which yielded primary frequencies that agreed well with the previous results (within +/-1 Hz, see Supplementary Material II for further details). Note, however, that the DMD analysis should be regarded as a mathematical tool that seeks the optimal approximation to the measurement data but does not always guarantee physically meaningful results, as it can be prone to various forms of experimental uncertainties. In contrast, the current study primarily relied on the phase-locked method (as described in Section 2.2) for cycle-averaged analysis.
3.4 Regime diagrams
The primary frequencies and modes of pulsating CH4-O2-CO2 flames, measured under various fuel flow rates, equivalence ratios, and dilution factors, are provided in Supplementary Material I. These results are also summarized in regime diagrams and frequency contour plots, as displayed in Figs. 8 and 9. Under the current experimental configuration, pulsating flame instability was observed only under fuel-rich conditions, specifically within the equivalence ratio range of 1.18 to 2.50. The range of unstable equivalence ratios narrows with increasing CO2 dilution (decreasing $\eta$), while its dependence on the CH4 flow rate is non-monotonic – first expanding with increasing CH4 flow rate, then peaking around 0.4 SLPM, and finally shrinking to a single point. This trend is consistent with the results reported previously in [27]. The highest CH4 flow rate at which flame pulsation was observed occurs near 1.32 SLPM, with the corresponding equivalence ratio and dilution factor being $\phi$ = 1.48 and $\eta$ = 1.00, respectively.
Figure 8: Regime diagrams of instability modes at selected CH4 mass flow rates. Lines indicate the observed stability boundaries.
Figure 9: Parametric dependence of the primary oscillation frequency of a pulsating CH4-O2-CO2 flame.
The single-mode oscillation regimes are generally narrower than those of multi-mode oscillations, except at low CH4 flow rates (0.25 SLPM or less), where heat losses to the burner are significant. At CH4 flow rates of 0.5 SLPM or higher, the single-mode regimes are confined to the vicinity of stability boundaries. The primary oscillation frequency is observed to increase with the mass flow rate of CH4 and decrease with the equivalence ratio, exhibiting a strong correlation with the flame’s thermal power. In contrast, the influence of the CO2 dilution factor on the primary oscillation frequency is less significant.
[…]
4 Conclusions
The pulsating instabilities of premixed CH4-O2 flames on a porous-plug burner were investigated at various levels of CO2 dilution, across a wide range of equivalence ratios, and flow rates. Time-resolved measurements of flame chemiluminescence and average temperature were obtained, from which the primary oscillation frequencies of the pulsating flames were determined using fast Fourier transform and harmonic power analysis. Additionally, spatiotemporally resolved measurements of gas temperature were conducted using spectrally modulated laser absorption diagnostics synchronized with high-speed chemiluminescence imaging.
From the results of the current measurements, a mode transition phenomenon in the pulsating flames was revealed through phase-locked analysis. Under fuel-rich conditions with relatively low heat release rates and flow rates, the flames exhibited quasi-periodic pulsation in a single mode. At elevated flow rates, flickering instability appeared and began to interfere with the thermo-diffusive pulsation, creating sidebands around the primary frequency. As the heat release rate increased, mode splitting occurred, and the flames eventually pulsated in multiple modes. The stability boundaries and regime diagrams of the flame oscillation modes were obtained, and representative examples of different modes were analyzed.
Although obtained from experiments on a specific porous-plug burner, the results of the current study – such as the overall trends of the stability boundary, primary frequency, and mode distributions – are expected to be broadly applicable to other CO2-diluted oxy-fuel combustion systems. They provide a basis for theoretical and numerical extrapolation toward applications relevant to practical combustion devices, where flame stability is a critical design consideration. While the current study primarily focused on methane flames, the underlying concepts can be easily extended to other fuels such as hydrogen and ammonia. In this sense, the present study should prove valuable for both fundamental research on flame dynamics and practical applications of CO2-moderated oxy-combustion.
Acknowledgments
This research is supported by the National Natural Science Foundation of China under Grants No. 92152108 and No. 12472278, and the Space Application System of China Manned Space Program under projects ”Nonlinear Dynamics of Flame Instability under Microgravity Conditions” and ”Ignition Mechanism of Premixed Near-Limit Flames under Microgravity Conditions”.
Supplementary Material
Supplementary Material I: Data summary of flame oscillation frequencies and modes under conditions explored in the current study. Supplementary Material II: Representative results of dynamic mode decomposition analysis.

### v3

@@
-A Parametric Study on the Pulsation of Burner-Stabilized CH4-O2 Flames Moderated by CO2 Addition – Frequencies, Modes and Regime Diagrams
+Pulsation of Burner-Stabilized CH4-O2 Flames Moderated by CO2 Addition – Frequencies, Modes and Regime Diagrams
 Abstract
@@
 Under fuel-rich conditions with relatively low heat release rates and low flow rates, the flames exhibited quasi-periodic single-mode oscillations.
-These oscillations were modulated by low-frequency flame flickering instabilities at elevated flow rates, which created sidebands around the primary oscillation frequency.
+At elevated flow rates, these oscillations were modulated by low-frequency flame flickering instabilities, which created sidebands around the primary oscillation frequency.
 At higher heat release rates, the flickering instability further triggered mode splitting, eventually leading to multi-mode oscillations.
@@
 Pulsating Instability; Oxy-Combustion; Porous-Plug Burner; Chemiluminescence; Laser Absorption Spectroscopy; Phase-Locked Analysis
-journal: Fuel
+journal: Combustion and Flame
 […]
@@
 Novelty and Significance
-This work presents, to the authors’ knowledge, the first systematic measurements of pulsating flame instabilities in burner-stabilized premixed methane-oxygen flames across a wide range of experimental conditions (equivalence ratios, dilution factors and flow rates).
-The combination of a porous-plug burner and oxygen-enriched combustion amplifies weak phenomena that commonly occur under extreme conditions of very high Lewis number, making them observable under moderate conditions that are easily accessible in a laboratory.
-Additionally, this work reports direct measurements of the spatiotemporal evolution of flame chemiluminescence and gas temperature during pulsation cycles.
-Based on phase-locked analysis of the data, this study has revealed an interesting mode-transition phenomenon of the flame pulsation.
-The results of this work should prove useful for both fundamental research on flame dynamics and practical applications of CO2-moderated oxy-combustion.
+This work presents, to the authors’ knowledge, the first systematic measurements of pulsating flame instabilities in burner-stabilized premixed methane–oxygen flames.
+Utilizing a custom-built miniature porous-plug burner and oxygen-enriched combustion, we obseverd the intrinsic flame pulsation typically occurring at very high Lewis numbers under moderate laboratory conditions.
+The spatiotemporal evolution of flame chemiluminescence and gas temperature was resolved throughout individual pulsation cycles, and a phase-locked analysis revealed an interesting mode-transition phenomenon of flame pulsation.
+The results of this work are useful for both fundamental research on flame dynamics and practical applications of CO2-moderated oxy-combustion.
 Author Contributions
@@
 An important type of such instability is the thermo–diffusive instability that arises from the imbalance between heat and mass transfer [7, 8, 9].
-Depending on the effective Lewis number (Le) of the reactive mixture, the flame can exhibit either spatial cellular structures (at Le $<$ 1) or temporal pulsation of the reaction front (at Le $>$ 1).
+Depending on the effective Lewis number (Le) of the reactive mixture, the flame can exhibit either spatial cellular structures (at Le << 1) or temporal pulsation of the reaction front (at Le >> 1).
 Although the former has been extensively investigated, especially in lean combustion systems of lightweight fuels such as hydrogen and methane [10, 11, 12, 13, 14, 15, 16, 17, 18], experimental studies on the latter remain relatively scarce.
@@
 A harmonic power analysis method was employed to extract the fundamental frequency of oscillation, $f_{osc}$, by exploiting the fact that the frequency content of a periodic physical event consists of harmonics that are phase-locked, whereas the frequency content of noise or random fluctuations is uncorrelated.
-Specifically, $f_{osc}$ was determined by solving a constrained optimization problem, namely maximizing the total power of harmonics within a search window $[f_{min},f_{max}]$ estimated from the NPSD result, as shown below.
+Specifically, $f_{osc}$ was determined by solving a constrained optimization problem, namely maximizing the total power of harmonics within a search window $f_{min},f_{max}]$ estimated from the NPSD result, as shown below.
 $f_{osc}=\underset{[f_{min},f_{max}]}{\operatorname{argmax}}\sum_{k=1}^{M}\left|\sum_{i=1}^{N}I(t_{i})e^{-2\pi jkft_{i}}\right|^{2}$ (3)
@@
 From the time sequence of $\alpha(t;y_{i})$, an absorbance spectrum $\alpha(\nu;y_{i})$ was obtained based on the laser wavenumber modulation profile $\nu(t)$ characterized prior to the flame experiment.
-A total of 12 absorption features in the spectral region of 2392.5 to 2397 $\rm cm^{-1}$ were analyzed, and the integrated absorbance over each selected transition feature (denoted by $[\nu_{k-},\nu_{k+}]$) was used in the axisymmetric reconstruction, i.e.,
-[…]
+A total of 12 absorption features in the spectral region of 2392.5 to 2397 $\rm cm^{-1}$ were analyzed, and the integrated absorbance over each selected transition feature (denoted by $\nu_{k-},\nu_{k+}]$) was used in the axisymmetric reconstruction, i.e.,
+$A_{k}(y_{i})=\int_{\nu_{k-}}^{\nu_{k+}}\alpha(\nu;y_{i})d\nu$ (5)
 To determine the spatial distribution of gas temperature, the entire spatial domain of interest was discretized using a number of sample points $r_{s}$ along the radial direction.
@@
 and
-[…]
+$A_{k}(y_{i})$ $=\int_{-\infty}^{+\infty}A^{0}_{k}\Big{(}\sqrt{x^{2}+y_{i}^{2}}\Big{)}dx\approx\sum_{s=1}^{\#pts.}w_{is}\cdot A_{k}^{0}(r_{s};T_{s},X_{s})$ (7)
 In the above equations, a constant flame pressure of $P$ = 1 atm was assumed, and the linestrength of each transition feature, $S_{k}^{0}$, was calculated as a function of the local gas temperature from the HITEMP spectroscopy model [46].

### v4

@@
-Pulsation of Burner-Stabilized CH4-O2 Flames Moderated by CO2 Addition – Frequencies, Modes and Regime Diagrams
+Pulsation of Burner-Stabilized CH4-O2 Flames Moderated by CO2 Addition
 Abstract
@@
 The flow rates of fuel, $\rm O_{2}$, and $\rm CO_{2}$ were precisely controlled by three Alicat MC series mass flow controllers, with typical uncertainties of 0.1%, 0.2%, and 0.5%, respectively.
-Multiple diagnostic methods were employed to investigate the flame oscillation dynamics, including: (1) a spatially and spectrally modulated multi-color laser absorption system for spatiotemporally resolved measurements of gas temperature and CO2 mole fraction; (2) an image-intensified high-speed camera for spatiotemporally resolved imaging of the line-of-sight integrated OH* chemiluminescence (as a proxy for the local heat release rate); and (3) a photomultiplier tube (PMT) for microsecond-resolved measurements of global chemiluminescence variations and transient phenomena.
-Figure 1: The current experimental setup: (a) schematic of the overall diagnostic system and burner; (b) detailed configuration of the porous-plug burner; (c) example of a burner-stabilized flame; (d) detailed view of the water cooling channel; (e) photograph of the static mixer.
+Multiple diagnostic methods were employed to investigate the flame oscillation dynamics, including: (1) a spatially and spectrally modulated multi-color laser absorption system for spatiotemporally resolved measurements of gas temperature (10 kHz, 2% accuracy); (2) an image-intensified high-speed camera for spatiotemporally resolved imaging of the line-of-sight integrated OH* chemiluminescence (10 kHz, as a proxy for the local heat release rate); (3) a photomultiplier tube (PMT) for microsecond-resolved measurements of global chemiluminescence variations and transient phenomena; (4) a Z-type high-speed schlieren system and (5) a pulsed OH-PLIF diagnostic for side-view measurements of the flames.
+Figure 1: The current experimental setup.
+(a) Configuration of the porous-plug burner.
+(b) Example of a burner-stabilized flame.
+(c) Detailed view of the water cooling channel.
+(d) Photograph of the static mixer.
+(e) Schematic of the flame chemiluminescence and the tunable diode laser diagnostics.
+(f) The side-view schlieren diagnostic.
+(g) The side-view OH-PLIF diagnostic system.
 The current laser absorption diagnostic used two distributed-feedback interband-cascade lasers (Nanoplus) to probe 12 absorption features of hot CO2 molecules near the $\nu_{3}$ bandhead at a scan rate of 10 kHz.
@@
 […]
+Side-view schlieren imaging of the flames was conducted using a Z-type configuration with a pair of concave mirrors, a 5-W LED, and a vertical knife edge located at the focal plane of the downstream mirror.
+The knife edge was slightly offset from the focus to enhance image contrast, enabling clear observation of minute density variations in the flow field during the experiment.
+Schlieren images were recorded with the Phantom v611 high-speed camera at a frame rate of 1 kHz.
+OH-PLIF measurements were performed using a nanosecond-pulsed tunable dye laser (LiOPStar-N, LIOP-TEC) at a repetition rate of 10 Hz.
+The laser wavelength was tuned to 282.997 nm, targeting the P(1.5) + Q(1.5) + R(2.5) transition cluster of the OH A–X (1,0) band, and the pulse energy was measured to be approximately 30 mJ.
+The laser beam was expanded into a horizontal light sheet using a set of cylindrical lenses, achieving a $2\sigma$ thickness of less than 0.3 mm across the burner surface.
+The OH-PLIF signal was collected with the same intensified camera used for OH* chemiluminescence, and spectrally filtered between 300–320 nm to suppress background flame emission.
+The camera was gated (with an exposure time of 10 $\mu s$) and synchronized to the laser pulses.
+During each OH-PLIF measurement, the PMT signal of OH* chemiluminescence was recorded simultaneously.
+The superposition of laser-induced fluorescence and laser scattering on OH* chemiluminescence produced local spikes in the PMT signal, enabling determination of the relative phase of each OH-PLIF image within the flame oscillation cycle.
 2.2 Frequency analysis and phase-locked cycle averaging of chemiluminescence signal
-For a pulsating flame, the frequency content of the total chemiluminescence signal was first analyzed using fast Fourier transform (FFT).
+The frequency content of the total chemiluminescence signal was first analyzed using fast Fourier transform (FFT).
 The results were expressed in terms of normalized power spectral density (NPSD), as shown in Eqn.
@@
 $\hat{S}(f_{k})=10\log_{10}S(f_{k})$ (2)
-A harmonic power analysis method was employed to extract the fundamental frequency of oscillation, $f_{osc}$, by exploiting the fact that the frequency content of a periodic physical event consists of harmonics that are phase-locked, whereas the frequency content of noise or random fluctuations is uncorrelated.
+The fundamental frequency of oscillation, $f_{osc}$, was extracted using a harmonic power analysis method developed in a previous study of the authors [52].
 Specifically, $f_{osc}$ was determined by solving a constrained optimization problem, namely maximizing the total power of harmonics within a search window $f_{min},f_{max}]$ estimated from the NPSD result, as shown below.
@@
 This optimization problem was solved numerically using a brute-force search algorithm, and to improve computational efficiency and robustness, the number of harmonics was truncated to M = 5.
-In the current study, this method was able to determine the primary oscillation frequency with a typical uncertainty of less than 0.1 Hz (defined by the frequency offset at 95% maximum total power), which was an order of magnitude better than the conventional FFT.
-Once the primary oscillation frequency was determined, the chemiluminescence images of the pulsating flames were further analyzed using phase-locked cycle averaging.
-The recorded image sequence was divided into sub-sequences corresponding to individual cycles of oscillation, and images captured at the same relative time within each cycle were averaged across different sub-sequences.
-The starting frame of each sub-sequence was fine-tuned through cross-correlation analysis, using the spatially integrated chemiluminescence signal as a reference, which effectively removed random jitters in the oscillation cycles.
-This procedure was executed iteratively until convergence was achieved, usually within 2-3 iterations.
-This phase-locked cycle averaging successfully removed random fluctuations and revealed the underlying physical trends consistent across cycles, without degrading the effective temporal resolution of the current measurements (100 $\mu s$).
+In the current study, this method was able to determine the primary oscillation frequency with a typical uncertainty of less than 0.1 Hz (defined by the frequency offset at 95% maximum total power).
+Once the primary oscillation frequency was determined, the chemiluminescence images of the pulsating flames were further analyzed using phase-locked cycle averaging, following the same procedure as explained in [52].
 2.3 Axisymmetric reconstruction of temperature fields from laser absorption measurement
@@
 For oscillating flames exhibiting axial symmetry, a reconstruction scheme using constrained optimization was developed to infer the spatial distribution of temperature from the absorbance data at different beam locations.
+Further details of the reconstruction scheme can be found in a previous study by the authors [58] and are therefore omitted here for brevity.
 Figure 2: Example measurement of a premixed CH4-O2 flame pulsating in a single axisymmetric mode under the conditions of $\dot{m}_{\rm CH_{4}}$ = 0.20 SLPM and $\dot{m}_{\rm O_{2}}$ = 0.30 SLPM.
@@
 (c) The Spatial distribution of gas temperature reconstructed from integrated absorbances under the assumption of axial symmetry.
-The line-of-sight absorbance $\alpha(t;y_{i})$ was determined at each beam location $y_{i}$ from the transmitted laser intensity $I(t;y_{i})$ and a reference intensity profile $I_{0}(t)$ measured in the absence of absorption, based on the Beer-Lambert relation:
-$\alpha(t;y_{i})$ $=-\ln\big{[}I(t;y_{i})/I_{0}(t)\big{]}$ (4)
-From the time sequence of $\alpha(t;y_{i})$, an absorbance spectrum $\alpha(\nu;y_{i})$ was obtained based on the laser wavenumber modulation profile $\nu(t)$ characterized prior to the flame experiment.
-A total of 12 absorption features in the spectral region of 2392.5 to 2397 $\rm cm^{-1}$ were analyzed, and the integrated absorbance over each selected transition feature (denoted by $\nu_{k-},\nu_{k+}]$) was used in the axisymmetric reconstruction, i.e.,
-$A_{k}(y_{i})=\int_{\nu_{k-}}^{\nu_{k+}}\alpha(\nu;y_{i})d\nu$ (5)
-To determine the spatial distribution of gas temperature, the entire spatial domain of interest was discretized using a number of sample points $r_{s}$ along the radial direction.
-These sample points were selected based on the effective spatial resolution of the current laser diagnostic (1 mm, as defined by the laser beam diameter).
-At each sample point, a local gas temperature $T_{s}$ and a CO2 mole fraction $X_{s}$ were assigned.
-The integrated absorbance of each absorption feature was evaluated first at each sample point $A^{0}_{k}(r_{s};T_{s},X_{s})$ and then summed along each line of sight (see Fig.
-2(b)), yielding:
-$A_{k}^{0}(r_{s};T_{s},X_{s})$ $=PX_{s}S_{k}^{0}(T_{s})$ (6)
-and
-$A_{k}(y_{i})$ $=\int_{-\infty}^{+\infty}A^{0}_{k}\Big{(}\sqrt{x^{2}+y_{i}^{2}}\Big{)}dx\approx\sum_{s=1}^{\#pts.}w_{is}\cdot A_{k}^{0}(r_{s};T_{s},X_{s})$ (7)
-In the above equations, a constant flame pressure of $P$ = 1 atm was assumed, and the linestrength of each transition feature, $S_{k}^{0}$, was calculated as a function of the local gas temperature from the HITEMP spectroscopy model [46].
-The line-of-sight integration was numerically evaluated using Gauss–Legendre quadrature, with linear interpolation between sample points.
-This approach effectively approximated the integral as a weighted sum of the integrated absorbances at selected sample points.
-The weight factor $w_{is}$ was a geometric constant independent of the laser wavenumber $\nu$ and the local gas properties $T_{s}$ and $X_{s}$.
-The integrated absorbances were assembled into two column vectors of $\boldsymbol{A^{0}(T,X)}=\{A^{0}_{k}(r_{s};T_{s},X_{s})\}$ and $\boldsymbol{A}=\{A_{k}(y_{i})\}$, with $\boldsymbol{T}$ and $\boldsymbol{X}$ representing $\{T_{i}\}$ and $\{X_{i}\}$, respectively.
-All weight factors $\{w_{is}\}$ were compiled into a constant matrix $\boldsymbol{W}$.
-The optimal values of $\boldsymbol{T}$ and $\boldsymbol{X}$ were determined by solving the following nonlinear optimization problem:
-$(\boldsymbol{T^{*},X^{*}})={\operatorname{argmin}}\big{\|}\boldsymbol{A-WA^{0}(T,X)}\big{\|}_{2}^{2}+\lambda_{T}^{2}\big{\|}\boldsymbol{LT}\big{\|}_{2}^{2}+\lambda_{X}^{2}\big{\|}\boldsymbol{LX}\big{\|}_{2}^{2}$ (8)
-The last two terms in Eqn.
-8 represent smoothness regularization, with $\boldsymbol{L}$ being the discrete Laplacian matrix and $\lambda_{T}$ and $\lambda_{X}$ serving as the regularization strength factors.
-The values of $\lambda_{T}$ and $\lambda_{X}$ were generally kept small to avoid blurring the reconstructed distributions, while being sufficiently large to ensure numerical stability and suppress random fluctuations; in the current study, they were set to $1\times 10^{-5}$ and $1\times 10^{-1}$, respectively.
-Eqn.
-8 was solved iteratively using the sequential quadratic programming method.
-The resulting spatial distributions of temperature and CO2 mole fraction are shown in Fig.
-4(c), with a typical measurement precision of 1%.
-Further discussion on the uncertainty analysis, as well as the choice of $\lambda_{T}$ and $\lambda_{X}$, can be found in a recent work by the authors [44].
 3 Results and Discussion
@@
 Figure 3: (a) Representative OH* NPSD spectra of pulsating flames at three different conditions.
-Top (blue): at $\phi$ = 1.72, $\eta$ = 1.00, the NPSD spectrum consists of equally spaced harmonics of a single fundamental frequency, indicating the presence of single-mode oscillation.
-Middle (green): at $\phi$ = 1.72, $\eta$ = 0.71, low-frequency flickering instability appears, which couples with the single-mode oscillation and creates sidebands around the harmonics.
-Bottom (red): at $\phi$ = 1.47, $\eta$ = 0.71, multiple fundamental oscillation frequencies (and their harmonics) are observed.
+Top (blue): $\phi$ = 1.72, $\eta$ = 1.00; middle (green): $\phi$ = 1.72, $\eta$ = 0.71; bottom (red): $\phi$ = 1.47, $\eta$ = 0.71.
 $f_{\text{OSC}}$: the primary oscillation frequency; $f_{\text{FI}}$: the flickering instability frequency.
@@
 At higher flow velocities, a second, low-frequency oscillation appeared around 10 Hz and created sidebands around the primary oscillation frequency, as shown in the middle panel of the figure.
-Across the experimental conditions explored in the current study, the frequency of this second oscillation was found to be relatively stable (9 - 12 Hz) and largely independent of the equivalence ratio and dilution level, suggesting that it must have originated from a mechanism different from the high-frequency thermo-diffusive pulsation.
-To further investigate the nature of this low-frequency oscillation, additional experiments were conducted using a CMOS camera to monitor the chemiluminescence signal of an undiluted, moderately fuel-rich flame from the side, as illustrated in Fig.
+Across the experimental conditions explored in the current study, the frequency of this second oscillation was found to be relatively stable (mostly between 9 and 12 Hz) and largely independent of the equivalence ratio and dilution level, suggesting that it must have originated from a mechanism different from the high-frequency thermo-diffusive pulsation.
+An example set of measurements for such low-frequency oscillation is shown in Fig.
 4.
-Under these conditions, the flame did not exhibit any high-frequency pulsation; only a low-frequency oscillation at a fundamental frequency of 11 Hz was observed.
+These measurements were conducted on an undiluted, moderately fuel-rich flame of equivalence ratio $\phi$ = 1.20.
+Under such conditions, the flame did not exhibit any high-frequency pulsation; only a low-frequency oscillation at a fundamental frequency of 11 Hz was observed.
 The results indicate that the low-frequency component corresponds to flickering-type flame dynamics that is distinctively different from flame front pulsation.
 For example, throughout a flame flickering cycle, the base of the flame remained largely stationary, whereas the tail of hot burnt gas downstream of the flame exhibited an oscillating motion of elongation and contraction along the axial direction.
-It turns out that the current observation of this rich premixed flame agrees well with classic studies on diffusion flames, in terms of both the overall phenomenon and the range of oscillation frequencies (e.g., [47, 48]).
-According to these studies, such low-frequency oscillations of the flame result from buoyancy-induced flow surrounding the flame.
-The velocity difference between the hot burnt gases and the cold ambient air creates an unstable shear layer, which rolls up into toroidal vortices and sheds periodically.
-The current study thus refers to this low-frequency oscillation as the flame flickering instability, with its frequency denoted as $f_{\rm FI}$.
-Figure 4: representative example on the low-frequency oscillation of a methane-oxygen flame at $\phi$ = 1.20, $\eta$ = 1.00.
-(a) the FFT spectrum of CH* normalized power spectral density (NPSD), (b) sequential side-view images of the flame.
+Figure 4: Example low-frequency oscillation of a methane-oxygen flame at $\phi$ = 1.20, $\eta$ = 1.00.
+(a) the FFT spectrum of OH* normalized power spectral density (NPSD), (b) sequential side-view images of the flame, (c) side-view schlieren images of the toroidal vortex formation and shredding.
+These observed low-frequency oscillations are consistent with the general frequency range of buoyancy-controlled flickering of diffusion flames reported in the early study by Buckmaster and Peters [7].
+The general dynamics of these oscillations agree qualitatively with the results of the planar visualization study by Chen et al.
+[12], in which the flickering instability is correlated with the periodic shredding of large toroidal vortices generated by buoyancy‑driven Kelvin–Helmholtz instability outside the luminous flame.
+The vortex-dynamical picture of flame flickering instability has been revisited by a multitude of studies under various fuel types, equivalence ratios, burner diameters, and gravity levels (for example, [44, 43, 54, 55]).
+A consensus has emerged that flickering is primarily caused by buoyant flow instability and toroidal vortices, at least at sufficiently small Reynolds numbers, and is distinct from flame extinction and re-ignition or direct modulation of the flame’s chemical reaction/heat release rate.
+In this sense, flickering instability is not restricted to diffusion flames but applies to premixed flames as well.
+The present study thus refers to this low-frequency oscillation as flame flickering instability, with its frequency denoted as $f_{\rm FI}$.
+Previous studies on the flickering frequency of buoyant diffusion flames (for example, [50, 4, 17, 11, 32, 21]) have suggested the Strouhal number ($St=f_{FI}D/U$, where $D$ is the burner diameter and $U$ is the unburnt gas velocity) correlates with the Froude number ($Fr=gD/U^{2}$) as $St\propto Fr^{-0.5}$.
+In this study, a total of 224 measurements were conducted on the oscillation frequencies of flickering flames without thermo-diffusive pulsations.
+The resulting non-dimensional frequency, expressed in terms of St, is plotted as a function of the inverse Froude number in Fig.
+5.
+The best numerical fit to the current data yields a scaling relation of $St\propto Fr^{-0.44}$, which agrees with the previous scaling of buoyancy-controlled flickering.
+Figure 5: Scaling of the measured non-dimensional flame flickering frequency (St) versus the inverse Froude number (1/Fr).
+Results are shown for purely flickering flames without thermo-diffusive pulsations.
+The present study also examined several other mechanisms that could potentially caused the low-frequency flame oscillations, including: (a) large-scale hydrodynamic shedding, (b) acoustic coupling, and (c) upstream forcing.
+None of these were found to be the actual cause.
+For hydrodynamic vortex shedding, the observed Strouhal number ($St$) varies significantly across different flame conditions, which differs markedly from the classical, nearly constant $St$ associated with vortex shedding from bluff bodies.
+Regarding acoustic coupling, the resonant acoustic frequency of the burner, estimated from the speed of sound and the effective cavity length inside the burner, is approximately 1250 Hz – two orders of magnitude higher than the observed flickering frequency.
+Upstream forcing was minimized by precisely controlling and continuously monitoring the mass flow rates throughout the experiments, and no upstream oscillation near 10 Hz was observed.
 At higher heat release rates (i.e., higher fuel flow rates and/or lower equivalence ratios), the sidebands around the primary oscillation frequency of the flame further split into two or more fundamental oscillation frequencies, each accompanied by its own harmonics, as shown in the bottom panel of Fig.
@@
 Eight representative frames of the cycle-averaged chemiluminescence images are shown in Fig.
-5(b) in pseudocolors, with red and blue representing high and low chemiluminescence signals, respectively.
+6(b) in pseudocolors, with red and blue representing high and low chemiluminescence signals, respectively.
 The mean intensity has been subtracted from each pixel to enhance contrast and highlight dynamic changes.
 From these images, the pulsating flame is observed to oscillate in a propagating ring mode, with concentric waves of local heat release rate perturbation originating from the edge and traveling toward the flame center, where they merge and terminate before the next oscillation cycle.
-It is worth noting that, although the measured images were line-of-sight integrated, no significant cancellation effect was expected in the top-view images because the flame front remained close to a geometrically thin sheet throughout the oscillation cycle.
-Besides, the flame motion occurred primarily in the vertical direction, as revealed by additional side-view measurement of a single-mode pulsating flame shown in Fig.
-6.
-Fig.
-5(c) illustrates the dynamic evolution of temperature distribution at a height above the burner (HAB) of 2.5 mm.
-A similar ring-type oscillation can be observed in the temperature distribution, but with a phase lag compared to the chemiluminescence signal.
+It is worth noting that, although the measured chemiluminescence images were line-of-sight integrated, no significant cancellation effect was expected in the top-view images because the flame front remained close to a geometrically thin sheet throughout the oscillation cycle.
+This was confirmed by side-view measurements of a single-mode pulsating flame, the results of which are shown in Fig.
+6(c).
+The dynamic evolution of the flame structure can also be visualized from OH-PLIF images, as shown in Fig.
+6(d).
+The PLIF measurements revealed additional details of the flame structure downstream of the flame front, which were obscured in the chemiluminescence imaging by the rapid quenching of electronically excited OH radicals (OH*) to their ground state.
+From the OH-PLIF measurements, an interesting phenomenon was observed: alternating regions of higher and lower OH signal were moving both horizontally and vertically within an oscillation cycle, suggesting that the flame oscillation mode followed a traveling-wave pattern.
+This was also confirmed by flame temperature measurements discussed below.
+The evolution of temperature distribution at a height above the burner (HAB) of 2.5 mm is illustrated in Fig.
+6(e).
+A similar ring-type oscillation can be observed.
 At the specific conditions of this experiment, the maximum local temperature is found to be approximately 2000 K, significantly lower than the adiabatic flame temperature ($\approx$ 3000 K) due to heat loss at the water-cooled burner surface.
 Fig.
-5(d) shows the evolution of line-of-sight-averaged temperature within one oscillation cycle.
+6(f) shows the evolution of line-of-sight-averaged temperature within one oscillation cycle.
 In the current study, a high level of measurement precision (on the order of $\pm$ 5 K) has been achieved through cycle averaging, thereby allowing small variations to be discerned.
-The average temperature is seen to oscillate between 1760 and 1790 K, with a phase lag of 100 - 150 degrees relative to the chemiluminescence signal.
+The average temperature is seen to oscillate between 1760 and 1790 K, with a phase delay of 100 - 150 degrees relative to the chemiluminescence signal.
 The overall 1-$\sigma$ uncertainties in these measurements are estimated to be 2 %, based on the uncertainty contributions from the spectral fitting (less than 0.1%), spectroscopy model (less than 1%), and small fluctuations in the laser intensity baseline (less than 1%).
-Fig.
-5(e) presents the vertical distribution of average temperature along the burner diameter, measured at different heights above the burner (HAB) and at various times during a single oscillation cycle.
-Its dynamic evolution reveals an interesting phenomenon that cannot be captured by the chemiluminescence imaging alone, namely, the existence of higher and lower average temperature regions and their vertical motion with time, which implies alternating puffs of hot and cool gas traveling in the vertical direction.
-Further investigation into the 2D/3D structure of the traveling flame fronts, especially in conjunction with velocity field measurement, may warrant future studies.
-Figure 5: Spatiotemporally resolved measurements of a single-mode pulsating flame ($\phi$ = 1.33, $\eta$ = 1.00) at different times in the oscillation cycle.
+Figure 6: Spatiotemporally resolved measurements of a single-mode pulsating flame ($\phi$ = 1.33, $\eta$ = 1.00) at different times in its oscillation cycle.
 (a) Spatially integrated chemiluminescence signals from the high-speed camera (black) and the PMT (blue), normalized and phase-locked averaged over 90 cycles.
-(b) Phase-locked averaged chemiluminescence images.
-(c) Dynamic evolution of the gas temperature distribution at HAB = 2.5 mm.
-(d) Evolution of the average temperature along the burner diameter.
+(b)-(c) Phase-locked averaged images of top-view and side-view chemiluminescence, respectively.
+(d) Instantaneous images of side-view OH-PLIF.
+(e) Dynamic evolution of the gas temperature distribution at HAB = 2.5 mm.
+(f) Evolution of the average temperature along the burner diameter.
 Error bars denote the standard deviation of the measurements.
-(e) Evolution of the vertical distribution of average gas temperature along the burner diameter within a single oscillation cycle.
-Figure 6: Side-view chemiluminescence images of a single-mode pulsating flame at $\phi$ = 2.00, $\eta$ = 1.00.
+(g) Evolution of the vertical distribution of average gas temperature along the burner diameter within a single oscillation cycle.
 Figure 7: Chemiluminescence signal illustrating the spatiotemporal modes of flame oscillation.
@@
 (a) Single-mode oscillation influenced by flickering instabilities at $\phi$ = 1.72, $\eta$ = 0.71.
-The dominant flame oscillation mode appears to be a fast centripetal traveling wave that is nearly axisymmetric.
-The flickering instabilities are manifested as a slow edge-mode perturbation that modulates the relative phase of the fast flame oscillation without significantly affecting its amplitude and spatial distribution.
 (b) Multi-mode oscillation at $\phi$ = 1.47, $\eta$ = 0.71.
-The integrated chemiluminescence intensity exhibits a clear beat pattern, indicating the presence of two fundamental oscillation frequencies that interfere with each other.
-The propagation directions and spatial patterns of flame oscillations during periods of constructive and destructive interference are drastically different.
+The vertical distribution of average temperature along the burner diameter, measured at different heights above the burner (HAB) and at various times during a single oscillation cycle, is presented in Fig.
+6(g).
+The existence of higher and lower average-temperature regions and their vertical motion with time correlates with the PLIF signal shown in Fig.
+6(d), indicating alternating puffs of hotter and cooler gases traveling in the vertical direction, with a phase delay similar to that of Fig.
+6(f).
+Such a phase delay suggests that the flame oscillation is not caused by heat loss to the burner surface alone, and there must be an additional mechanism – such as the differential transport between mass and heat – that interacts with the heat loss and jointly drives the oscillation cycle.
+Further investigation into the 2D/3D structure of the traveling flame fronts and the detailed mechanism of flame oscillation, especially in conjunction with velocity-field measurement, is warranted for future studies.
 3.3 Multi-mode oscillation
-Two more complex cases of pulsating flames are presented in Fig.
-7, including: (a) single-mode oscillations coupled with flickering instabilities, and (b) multi-mode oscillations.
-In the former case, the dominant flame oscillation mode remains a centripetal traveling wave, but flickering instabilities introduce a slow edge-mode perturbation that modulates its relative phase without significantly affecting its amplitude and spatial distribution.
-In the latter case, the amplitude of the flame oscillation varies significantly over time.
+Fig.
+7 shows two representative cases of pulsating flames exhibiting more complex dynamics, including: (a) single-mode oscillations coupled with flickering instabilities, and (b) multi-mode oscillations.
+In the former case, the dominant flame oscillation mode remains a ring-shaped traveling wave, but flickering instabilities introduce a slow edge-mode perturbation that modulates its relative phase without significantly affecting its amplitude and spatial distribution.
+In the latter case, the amplitude of the flame oscillation varies more drastically over time.
 For example, in a representative measurement shown in Fig.
 7(b), the integrated chemiluminescence intensity exhibits a clear beat pattern, indicating the presence of two fundamental oscillation frequencies interfering with each other.
-During the time interval of constructive interference, the spatial pattern of the flame oscillation remains axisymmetric, while the propagation of the heat release rate perturbation becomes bidirectional – waves originate from a ring zone and travel toward both the center and the edge of the burner.
-In contrast, during the time interval of destructive interference, the flame oscillation pattern appears drastically different: perturbation waves travel along two separate loops near the top and bottom edges of the burner and meet along the diameter.
-More complex interactions among three or more modes have also been observed at much higher heat release rates.
+During their constructive interference, the spatial pattern of the flame oscillation appears axisymmetric – perturbation waves emerge from a ring zone and travel toward both the center and the edge of the burner.
+In contrast, during their destructive interference, the flame oscillation pattern appears remarkably different: perturbation waves travel along two separate loops near the top and bottom edges of the burner and converge along the diameter.
+More complex interactions among three or more modes have also appeared at conditions of higher heat release rates.
+In the present study, a flame was classified as stable, flickering, single-mode, or multi-mode pulsating based on three parameters: (1) the dominant/strongest frequency, $f_{osc}$, defined by Eqn.3; (2) the signal-to-noise ratio (SNR) at $f_{osc}$, evaluated in relative units of dB with respect to the average noise floor ($\hat{S}_{noise}$ = -80 dB/Hz) across 100 - 500,000 Hz, i.e., SNR $\hat{S}(f_{osc})-\hat{S}_{noise}$; and (3) the relative spectral power ($R$) of the second-strongest pulsating frequency ($f_{osc,2}$) with respect to the strongest frequency ($f_{osc}$), i.e., $R=\hat{S}(f_{osc,2})-\hat{S}(f_{osc})$.
+First, a flame was classified as oscillating if SNR >> 30 dB; otherwise, it was considered stable.
+It is worth noting that the main conclusions of the present study were not sensitive to the specific choice of this threshold, as the qualitative results remained largely unchanged when its value was varied within $\pm$ 10 dB.
+Then, an oscillating flame was classified as flickering if $f_{osc}<$ 20 Hz; otherwise, it was classified as pulsating.
+As illustrated in Fig.
+8, this classification between the flickering mode and the pulsating mode is self-explanatory.
+Finally, a pulsating flame was classified as multi-mode if $R>$ -10 dB (representing that the spectral power of the second strongest mode is higher than 10% of the strongest mode); otherwise, it was classified as single-mode.
+Note that, in principle, the transition between single-mode and multi-mode should be continuous and does not necessarily involve a definite boundary (in a sense, the value of $R$ itself can be regarded as an objective metric of mode purity).
+However, to facilitate discussion of the current results, a -10 dB criterion was imposed.
+As demonstrated later in Fig.
+9, this threshold value closely captures the steepest slope of $R$ with respect to $\phi$, representing a reasonably good separation of modes.
+[…]
+Figure 8: Joint distribution of the observed $f_{osc}$ and SNR from the current experiments.
 Dynamic Mode Decomposition (DMD) analysis was also performed on selected OH* image sequences, which yielded primary frequencies that agreed well with the previous results (within +/-1 Hz, see Supplementary Material II for further details).
@@
 These results are also summarized in regime diagrams and frequency contour plots, as displayed in Figs.
-8 and 9.
+9 and 10.
 Under the current experimental configuration, pulsating flame instability was observed only under fuel-rich conditions, specifically within the equivalence ratio range of 1.18 to 2.50.
@@
 The highest CH4 flow rate at which flame pulsation was observed occurs near 1.32 SLPM, with the corresponding equivalence ratio and dilution factor being $\phi$ = 1.48 and $\eta$ = 1.00, respectively.
-Figure 8: Regime diagrams of instability modes at selected CH4 mass flow rates. Lines indicate the observed stability boundaries.
+Figure 9: Regime diagrams of instability modes at selected CH4 mass flow rates.
+Solid lines: the observed stability boundaries.
+Dash lines: the boundary between single-mode and multi-mode oscillations defined by R=10R=-10 dB.
 Figure 10: Parametric dependence of the primary oscillation frequency of a pulsating CH4-O2-CO2 flame.
@@
 In contrast, the influence of the CO2 dilution factor on the primary oscillation frequency is less significant.
-[…]
+The dependence of the effective Lewis number $Le$ on the dilution factor $\eta$ and the equivalence ratio $\phi$ was evaluated from 1-D freely-propagating adiabatic flame simulations performed with Cantera [15] using the DRM19 mechanism [23].
+The results are shown in Fig.
+11.
+Across the range of conditions explored in the current study, $Le$ varies between 0.96 and 1.09.
+Figure 11: Effective Lewis number ($Le$) as a function of equivalence ratio ($\phi$) and dilution factor ($\eta$), calculated using the DRM19 mechanism [23].
+A non-monotonic dependence of $Le$ on CO2 addition is observed, arising from competing effects of mass and thermal diffusion.
+On the one hand, CO2 addition reduces the mixture-averaged species diffusivity and slows mass transport; on the other hand, it increases the mixture’s specific heat capacity and reduces the overall thermal diffusivity, especially at high dilution.
+The reduction in mass diffusivity dominates at low CO2 dilution ($0.8\leq\eta\leq 1.0$), producing a negative dependence of $Le$ on $\eta$; whereas at relatively high CO2 dilution ($\eta<0.8$), the large CO2 heat capacity becomes the dominant factor and yields a positive dependence of $Le$ on $\eta$.
+At conditions of relatively high $Le$ (e.g., $Le>1.08$), the flame pulsation was multi-mode.
+Throughout the current study, no flame pulsation was observed at conditions of $Le<1$.
 4 Conclusions
-The pulsating instabilities of premixed CH4-O2 flames on a porous-plug burner were investigated at various levels of CO2 dilution, across a wide range of equivalence ratios, and flow rates.
-Time-resolved measurements of flame chemiluminescence and average temperature were obtained, from which the primary oscillation frequencies of the pulsating flames were determined using fast Fourier transform and harmonic power analysis.
-Additionally, spatiotemporally resolved measurements of gas temperature were conducted using spectrally modulated laser absorption diagnostics synchronized with high-speed chemiluminescence imaging.
-From the results of the current measurements, a mode transition phenomenon in the pulsating flames was revealed through phase-locked analysis.
-Under fuel-rich conditions with relatively low heat release rates and flow rates, the flames exhibited quasi-periodic pulsation in a single mode.
-At elevated flow rates, flickering instability appeared and began to interfere with the thermo-diffusive pulsation, creating sidebands around the primary frequency.
-As the heat release rate increased, mode splitting occurred, and the flames eventually pulsated in multiple modes.
-The stability boundaries and regime diagrams of the flame oscillation modes were obtained, and representative examples of different modes were analyzed.
-Although obtained from experiments on a specific porous-plug burner, the results of the current study – such as the overall trends of the stability boundary, primary frequency, and mode distributions – are expected to be broadly applicable to other CO2-diluted oxy-fuel combustion systems.
-They provide a basis for theoretical and numerical extrapolation toward applications relevant to practical combustion devices, where flame stability is a critical design consideration.
-While the current study primarily focused on methane flames, the underlying concepts can be easily extended to other fuels such as hydrogen and ammonia.
-In this sense, the present study should prove valuable for both fundamental research on flame dynamics and practical applications of CO2-moderated oxy-combustion.
+The pulsating instabilities of premixed CH4-O2 flames on a porous-plug burner were investigated at various levels of CO2 dilution, equivalence ratios, and flow rates.
+Heat loss from the flat burner surface amplified the thermo-diffusive instability of these flames and lead to pulsations at near-unity Lewis numbers where freely propagating flames would be stable.
+In the present study, different modes of flame pulsation were observed.
+Under fuel-rich conditions with relatively low heat release and flow rates, the flames exhibited quasi-periodic pulsation in a single mode.
+At elevated flow rates, buoyancy-controlled, low-frequency flickering instability appeared and began to interfere with the thermo-diffusive pulsation, creating sidebands around the primary frequency.
+At elevated heat release rates, nonlinear interaction between flame flicker and single-mode pulsation further triggered mode splitting, and multi-mode pulsation emerged.
+The stability boundaries and regime diagrams of the flame oscillation modes were established, and representative measurements of the spatiotemporal variations in pulsating flames were analyzed.
+The general trends of the stability boundary and mode distributions are generalizable to other CO2-diluted oxy-fuel combustion systems.
+While this study focuses on methane flames, the underlying concepts can be extended to other fuels, such as hydrogen and hydrogen-blended fuels.
+In this sense, the present study should prove valuable for both fundamental research on flame dynamics and practical applications of oxy-combustion.
 Acknowledgments
-This research is supported by the National Natural Science Foundation of China under Grants No.
+This research is supported by the National Key Research and Development Program of China under Grant No.
+2025YFF0511801, the National Natural Science Foundation of China under Grants No.
 12472278 and No.
-12472278, and the Space Application System of China Manned Space Program under projects ”Nonlinear Dynamics of Flame Instability under Microgravity Conditions” and ”Ignition Mechanism of Premixed Near-Limit Flames under Microgravity Conditions”.
+92152108, the Space Application System of China Manned Space Program under projects ”Nonlinear Dynamics of Flame Instability under Microgravity Conditions” and ”Ignition Mechanism of Premixed Near-Limit Flames under Microgravity Conditions”, and the Open Research Program of National Key Laboratory of Fundamental Algorithms and Models for Engineering Simulation.
 Supplementary Material